# CricaVPR: Cross-image Correlation-aware Representation Learning for Visual Place Recognition

### v1

@@ 14 Datasets Details
Pitts30k [57] is derived from Google Street View panoramas with GPS labels. It consists of images from 24 different viewpoints for each place in urban scenes, exhibiting significant viewpoint variations, moderate condition variations, and a small number of dynamic objects. For our experiments, we use the Pitts30k test set, which comprises 10k database images and 6,816 query images.
@@
MSLS (Mapillary Street-Level Sequences) [64] is a large-scale VPR dataset that encompasses more than 1.6 million images captured in urban, suburban, and natural environments across 30 cities spanning six continents. This dataset provides GPS coordinates and compass angles for each image, and shows various changes caused by illumination, weather, season, viewpoint, dynamic objects, and so on. It is divided into three sets: training, public validation (MSLS-val), and withheld test (MSLS-challenge). To ensure comprehensive evaluation, we assess the model on both the MSLS-val and MSLS-challenge sets, as done in previous works [26, 61].
@@
AmsterTime [69] contains more than a thousand query-reference image pairs captured from Amsterdam. Each pair consists of a grayscale historical image as the query and a contemporary image from the same place (identified by human experts) as the reference. The dataset involves diverse domain variations in viewpoints, modalities (RGB vs grayscale), very long-term time spans, etc., which makes it quite difficult for VPR.
@@ 15 Compared Methods Details
Besides, the results on the three challenging datasets (Nordland, AmsterTime, and SVOX) are directly referenced from the EigenPlaces paper [10]. These results are basically consistent with what we reproduced.
@@ 16 Additional Qualitative Results and Failure Cases
In the main paper, we have presented a small number of qualitative results to show the robustness of our approach in challenging scenarios. In this section, we add more examples to vividly demonstrate the performance of VPR methods. Fig. 8, Fig. 9, and Fig. 10 show examples on Pitts30k, Tokyo24/7, and MSLS-val, respectively. These examples demonstrate that our method is more robust against variations in conditions and viewpoints, as well as perceptual aliasing, than previous methods. Fig. 11 and Fig. 12 show examples on AmsterTime and Nordland, which demonstrate that our method can correctly recognize place images over long time spans and under extreme environments in general. However, our method also produces erroneous results in a few cases (the last examples of these two figures) when images from different places are very similar, especially when lacking discriminative landmarks. This is a focus to improve our method in the future.

### v2

@@ 15 Datasets Details
Pitts30k [58] is derived from Google Street View panoramas with GPS labels. It consists of images from 24 different viewpoints for each place in urban scenes, exhibiting significant viewpoint variations, moderate condition variations, and a small number of dynamic objects. Pitts30k is a subset of Pitts250k (but harder than Pitts250k for most methods). In our experiments, we mainly use the Pitts30k test set.
@@
MSLS (Mapillary Street-Level Sequences) [65] is a large-scale VPR dataset that encompasses more than 1.6 million images captured in urban, suburban, and natural environments across 30 cities spanning six continents. This dataset provides GPS coordinates and compass angles for each image, and shows various changes caused by illumination, weather, season, viewpoint, dynamic objects, and so on. It is divided into three sets: training, public validation (MSLS-val), and withheld test (MSLS-challenge). To ensure comprehensive evaluation, we assess the model on both the MSLS-val and MSLS-challenge sets, as done in previous works [26, 43, 44].
@@
AmsterTime [70] contains more than a thousand query-reference image pairs captured from Amsterdam. Each pair consists of a grayscale historical image as the query and a contemporary image from the same place (identified by human experts) as the reference. The dataset involves very long-term time spans, and diverse domain variations in viewpoints, modalities (RGB vs grayscale), etc., which makes it quite difficult for VPR.
@@ 16 Compared Methods Details
Besides, the results on the three challenging datasets (Nordland, AmsterTime, and SVOX) are directly referenced from the EigenPlaces paper [10]. These results are basically consistent with what we have reproduced.
@@ 17 Additional Qualitative Results and Failure Cases
In the main paper, we have presented a small number of qualitative results to show the robustness of our approach in challenging scenarios. In this section, we add more examples to vividly demonstrate the performance of VPR methods. Fig. 8, Fig. 9, and Fig. 10 show examples on Pitts30k, Tokyo24/7, and MSLS-val, respectively. These examples demonstrate that our method is more robust against variations in conditions and viewpoints, as well as perceptual aliasing, than previous methods. Fig. 11 and Fig. 12 show examples on AmsterTime and Nordland, which demonstrate that our method can correctly recognize place images over long time spans and under extreme environments in general. However, our method also produces erroneous results in a few cases (the last examples of these two figures) when images from different places are very similar, especially when lacking discriminative landmarks.
18 Limitations
In addition to the failure case mentioned in the previous section, our approach has two limitations. First, although our approach achieves excellent results with the 512-dim compact feature on Pitts30k, it does not perform well on datasets with severe condition changes (e.g., Tokyo247, MSLS) when the descriptor dimension is reduced to very low. Secondly, setting the inference batch size to 1 will render our cross-image encoder ineffective, resulting in a gap between training and testing, and thus not achieving optimal performance. These are the focal points for future improvement of our method.